OpenStudy (tdskitz):

Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (jonathan34):

ok

OpenStudy (princeharryyy):

|x+2|=14-16 |x+2| = -2 case I>> x+2 = -2 x=-4 case 2 >> x+2 = +2 x=0

OpenStudy (tdskitz):

thx

OpenStudy (princeharryyy):

:)